Let f(x)=logx, x>0
Write down the value of
f(a)
f(1)
f(a^4)

Answer by stanbon(75887)

Let f(x)=logx, x>0
Write down the value of
f(a) = log(a)
------
f(1) = log(1) = 0
------
f(a^4) = log(a^4) = 4*log(a)
-------------
Cheers,
Stan H.
------------